The total age of family of 4is 68 4 years ago it was 53 how old is the youngest
Found 2 solutions by solver91311, Edwin McCravy:
Answer by solver91311(24713)



If the youngest was 4 or more, the sum would be 16 less. If the youngest was 2, the sum would be 14 less and if the youngest were 1, the sum would be 13 less. But the sum is 15 less, so the youngest has to be 3.

John

My calculator said it, I believe it, that settles it

Answer by Edwin McCravy(20055)



Maybe we should use some algebra in there somewhere instead
of using trial and error.

Normally we would expect that if the sum of the 4 people's
ages is now 68, then four years ago each would have been 4 
years younger, and the sum would have been 4*4 or 16 less.
But 16 less than 68 is 52, not 53. 

So the youngest child must not have been born 4 years ago.

So 4 years ago there were only 3 members of the family

Let the sum of the ages of the 3 older members of the family be x.
Let the age of the youngest child who wasn't born 4 years ago be y.

Then x + y = 68 and since 4 years ago there were only 3 members of
the family the sum would have been only 4*3 or 12 less. So

x-12 = 53
x = 65

And since x + y = 68, substituting x = 65
         65 + y = 68
              y = 3

The youngest child is 3.

Edwin